Naruko885:

A standard showerhead in Jen's house dispenses 5 gallons of water per minute. Jen changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Jen save each day with the change in showerheads if she uses the shower for 8 minutes each day? 16 gallons 17 gallons 32 gallons 37 gallons

Naruko885:

@Hanna sdo you have any idea?

Zarkon:

looks like the new shower head only gives 3 gallons per minute. That is a saving of 2 gallons per minute (assuming y is in gallons)

psirockin2:

y=5(8)=40 y=3(8)=24 it saves 16 gallons when you switch shower-heads

Zarkon:

or 2 gallons saved per minute times 8 minutes... 2*8=16